### v2

# Causal Data Fusion for Panel Data without Pre-Intervention Period

Abstract
Traditional panel data causal inference frameworks, such as difference-in-differences and synthetic control methods, rely on pre-intervention data to estimate counterfactuals. However, such data may not be available in real-world settings when interventions are implemented in response to sudden events, such as public health crises or epidemiological shocks. In this paper, we introduce two data fusion methods for causal inference from panel data in scenarios where pre-intervention data is unavailable. These methods leverage auxiliary reference domains with related panel data to estimate causal effects in the target domain, overcoming the limitations imposed by the absence of pre-intervention data. We show the efficacy of these methods by obtaining converging bounds on the bias as well as through a simulation study. Our proposed methodology renders causal inference feasible in urgent and data-constrained environments where the assumptions of the existing causal inference frameworks are not met. As an application of the proposed methodology, we study the causal effect of the community organization activity on the COVID-19 vaccination rate among the Hispanic sub-population in the city of Chelsea, Massachusetts.
Keywords: Causal Inference; Data Fusion; Panel Data; Equi-Confounding; Synthetic Control
1 Introduction
In many real-world applications, a treatment or intervention is applied to a target unit in the system, and the researcher is interested in assessing the causal effect of this intervention by evaluating the counterfactual behavior of the target unit in the absence of the intervention. This setting is challenging because, unlike conventional causal inference scenarios, there is only one treated unit, and there may be unobserved variables in the system which have confounded the treatment-outcome relationship.
The difference-in-difference causal inference framework (Card,, 1990; Angrist and Pischke,, 2008) and its significant extension, the synthetic control framework (Abadie and Gardeazabal,, 2003; Abadie et al.,, 2010, 2015), are well-known panel data methods commonly used in such settings. These methods require access to pre-intervention data from the target unit to match it with units (or combinations of units) from the control group. However, in some applications, such pre-intervention data may not be available. For example, in the case of a sudden shock or emergency, immediate action may be taken by the agencies, and hence, no data will be available from the unit during the emergency before the action was taken. For instance, during the COVID-19 pandemic, interventions were implemented almost instantaneously to curb the spread of the virus. As an example, when COVID-19 vaccines first became available in Massachusetts, vaccine doses were accessible only to English-fluent people who were able to use online appointment scheduling sites and who could drive to mass vaccination centers. For many low-income residents who lacked access to information in their preferred language, to a computer or to a car, COVID-19 vaccines remained out of reach, and efforts of a community-based organization in one city, the city of Chelsea, became exemplary for improving vaccine access for their community members. As the organization was instituted (which is the intervention in the system) from almost the first day, these unprecedented circumstances left researchers with no pre-intervention data from the target units to analyze the impact of the interventions. This lack of pre-intervention data hampers the ability to draw robust conclusions about the effectiveness of the interventions and inform future policy decisions.
To address this challenge, in this work, we propose two data fusion-based causal inference methods for panel data with no pre-intervention period. We consider settings where besides the target domain in which we aim to assess the causal effect, we have access to panel data from the units in an auxiliary reference domain. The reference domain can be on time periods before, concurrent or after the time periods in the target domain. Moreover, the observed covariates in the reference domain can be different from those in the target domain. Our first proposed method, called equi-confounding data fusion, is an extension of the ideas of the difference-in-differences approach, and our second proposed method, called synthetic control data fusion, is an extension of the ideas of the celebrated synthetic control method. As mentioned earlier, both the standard difference-in-differences and synthetic control methods require pre-intervention data. Therefore, our contribution is to formalize sufficient conditions regarding the relation between the data in the reference and target domains that can render the identification and estimation of the causal effect using integration (fusion) of the data from the two domains feasible. These conditions essentially allow us to anchor the counterfactual outcome of the unit of interest in the target domain at that in the reference domain. A challenge introduced by having more than one domain is that extra caution should be taken in matching the target unit with the control units, as naive extension of the existing matching methods may lead to unsatisfactory match in one or both of the domains. We address this issue by introducing a novel data-driven constrained optimization method for designing matches in our proposed synthetic control data fusion approach. We show the efficacy of these methods by obtaining bounds on the absolute biases, and provided conditions under which the bounds converge to zero. The use of data fusion for identification and estimation of causal parameters in settings where data from one domain is not sufficient for effective evaluation of causal parameters has gained significant attention in the recent years (Dahabreh et al.,, 2020; Degtiar and Rose,, 2023; Colnet et al.,, 2024). However, to the best of our knowledge, we are the first to study this approach for the challenging setting of panel data with single target unit where latent confounding variables may be present.
The rest of the paper is organized as follows. We provide a formal description of the problem in Section 2 and lay out the basic assumptions of the setting under study. We present our proposed equi-confounding data fusion and synthetic control data fusion in Sections 3 and 4, respectively. We evaluate our proposed approaches on synthetic data in Section 5. We apply our proposed methodologies to the application of COVID-19 vaccination rate evaluation in Section 6. Our concluding remarks are provided in Section 7.
2 Problem Description
We consider a setting with $J+1$ units observed over $S$ time periods. $A_{i,s}\in\{0,1\}$ and $Y_{i,s}\in\mathbb{R}$ denote the treatment and outcome random variables for unit $i$ at time period $s$, respectively, and $X_{i}\in\mathbb{R}^{d_{t}}$ denotes the observe covariates for unit $11\leq i\leq J+$ 1. We consider the case where unit $i=1$ receives the treatment over all time periods, and the rest of the $J$ units are control units, i.e., they never receive treatment. Formally,
$A_{i,s}=\begin{cases}1&\text{if }i=1,\text{ and }1\leq s\leq S,\\ 0&\text{if }2\leq i\leq J+1\text{ and }1\leq s\leq S.\end{cases}$
We refer to unit $i=1$ as the target unit. An example of this setting is the study of effect of community organizations on COVID-19 vaccination rate a sub-population in cities, mentioned in Section 1. Here, the target city had a strong community organization with a large team of community health workers and the ability to provide wraparound services for fundamental social determinants of health from the beginning of the rollout of the vaccines, and the rest of the cities under consideration did not have such community organizations.
We use the potential outcomes framework (Rubin,, 1974) and denote the potential outcome variable of $Y_{i,s}$, had the treatment variables been set to value $A_{i,s}=a$, for all $s$ (possibly contrary to the fact) by $Y_{i,s}^{(a)}$, where $a\in\{0,1\}$. Our parameter of interest is the effect of the treatment on the treated (ETT) averaged over the $S$ treatment time periods, that is,
$\psi_{0}=\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}[Y_{i,s}^{(1)}-Y_{i,s}^{(0)}\mid A_{i,s}=1\text{ for all }s].$
Since unit $i=1$ is the only unit with $A_{i,s}=1$, the parameter of interest, $\psi_{0}$, can be written as
$\psi_{0}=\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}[Y_{1,s}^{(1)}-Y_{1,s}^{(0)}].$ (1)
We assume the consistency assumption (Hernán and Robins,, 2020) for the variables, formally stated as follows.
Assumption 1 (Consistency).
For all $11\leq i\leq J+$ 1 and $1\leq s\leq S$, we have $Y_{i,s}=Y_{i,s}^{(a)}$ if $A_{i,s}=a$ for all $s$.
Due to the consistency assumption, $\psi_{0}$ can be written as $\psi_{0}=\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}[Y_{1,s}-Y_{1,s}^{(0)}]$. Therefore, the challenge in evaluating the causal effect is assessing the potential outcome mean of the target unit under control. Similar setting has been considered in the literature of causal inference in the difference-in-differences framework, synthetic control framework, and their variations (Arkhangelsky and Imbens,, 2024). However, in all those approaches, it is assumed that we also observed the target unit before the implementation of the treatment, which is not the case in our setting. Therefore, those approaches cannot be applied to our setting.
In our setting of interest, we also have access to data from a secondary domain from the same units. We refer to the primary domain, in which we are interested to evaluate the causal parameter as the target domain, and the secondary domain as the reference domain. In the reference domain, units are observed over $T$ time periods. We denote the observed covariates of the units in this domain by $Z_{i}\in\mathbb{R}^{d_{r}}$, $11\leq i\leq J+$ 1. Note that we do not necessarily observe the same covariates for the units in the target and the reference domains and $d_{r}$ is not necessarily equal to $d_{t}$. We also observe a certain outcome variable $F_{i,t}\in\mathbb{R}$ for $11\leq i\leq J+$ 1 and $1\leq t\leq T$ in the reference domain.
Our goal is to formalize conditions regarding the relevance of the outcome variables $F_{i,t}$ in the reference domain and $Y_{i,s}$ in the target domain, under which the observations in the reference domain can be leveraged for evaluating the causal parameter of interest in the target domain. To this end, we propose two data fusion approaches, named equi-confounding data fusion and synthetic control data fusion, which are described in Sections 3 and 4, respectively.
3 Equi-Confounding Data Fusion
3.1 Linear Equi-Confounding
Our first proposed approach is an extension of the ideas of negative controls and the difference-in-differences causal inference frameworks to our data fusion setup. Similar extension appeared in (Ghassami et al.,, 2022) for evaluating the long-term causal effect of treatments via data fusion.
Let $Y^{(0)}_{i}:=\frac{1}{S}\sum_{s=1}^{S}Y^{(0)}_{i,s}$, $Y_{i}:=\frac{1}{S}\sum_{s=1}^{S}Y_{i,s}$, and $F_{i}:=\frac{1}{T}\sum_{T=1}^{T}F_{i,t}$. The main assumption of our approach is the following.
Assumption 2 (Linear equi-confounding).
The potential outcome variables in the target domain and the outcome variables in the reference domain satisfy the following.
$\mathbb{E}[Y_{1}^{(0)}]-\frac{1}{J}\sum_{i=2}^{J+1}\mathbb{E}[Y_{i}^{(0)}]=\mathbb{E}[F_{1}]-\frac{1}{J}\sum_{i=2}^{J+1}\mathbb{E}[F_{i}].$ (2)
The left hand side in Equation (2) is based on the variables in the target domain and the right hand side is based on the variables in the reference domain. Hence, Assumption 2 specifies the connection between the outcomes in the target domain and the outcomes in the reference domain, and clarifies how the information in the reference domain can be transferred to the target domain. The quantity $\mathbb{E}[Y_{1}^{(0)}]-\frac{1}{J}\sum_{i=2}^{J+1}\mathbb{E}[Y_{i}^{(0)}]$ measures the confounding effect on the potential outcome $Y^{(0)}_{1}$. Therefore, Assumption 2 requires that such effect for $Y^{(0)}_{1}$ and $F_{1}$ are equal; hence, the name equi-confounding. The assumption is inspired by the additive equi-confounding bias assumption in the identification approach using negative outcome controls (Lipsitch et al.,, 2010; Tchetgen Tchetgen,, 2014; Sofer et al.,, 2016), and is similar in flavor to the parallel trends assumption in the difference-in-differences framework (Card,, 1990; Angrist and Pischke,, 2008).
To gain intuition about Assumption 2, suppose the intervention in the target domain does not have any causal effect on the outcome variables in the reference domain. For instance, in the vaccine rollout application, this may be satisfied, if we consider the outcome variable in the reference domain as the vaccination rate for another sub-population, for example, the Black sub-population, which is not able to be equally targeted by the community organization. Another example is the influenza vaccination rate for the same sub-population in the case that the community organization did not campaign for the influenza vaccines. Then, Assumption 2 requires that the discrepancy in the mean of the potential outcomes in the absence of any community organization activities in the target domain mirrors that in the reference domain. For instance, the difference in the mean of COVID-19 vaccination rate for the Black sub-population between the target unit and the average of the rest of the units is the same as the same quantity for COVID-19 vaccination rate for the Hispanic sub-population if the target unit (i.e., the city of Chelsea) did not receive any influence from the community organization.
Below, we demonstrate an example of a model in which Assumption 2 is satisfied.
Example 1.
Assumption 2 is satisfied if the data is generated from the following structural equations.
$Y_{i,s}=\varrho_{s}+f(G_{i},X_{i})+\alpha_{i,s}A_{i,s}+\varepsilon_{i,s},$
$F_{i,t}=\rho_{t}+g(G_{i},Z_{i})+\epsilon_{i,t},$
with $X_{i}\overset{i.i.d.}{\sim}P_{X}$ and $Z_{i}\overset{i.i.d.}{\sim}P_{Z}$, for some distributions $P_{X}$ and $P_{Z}$, and with $\mathbb{E}[f(1,X_{i})-f(0,X_{i})]=\mathbb{E}[g(1,Z_{i})-g(0,Z_{i})]$, where $\varepsilon_{i,s}$’s and $\epsilon_{i,t}$’s are mean-zero independent noise variables, and $G_{i}$ is a target unit indicator variable which is equal to 1 if $i=1$ and is equal to zero, otherwise.
Consider the following estimator for the causal parameter $\psi_{0}$.
$\hat{\psi}^{eq1}=\left\{Y_{1}-F_{1}\right\}-\frac{1}{J}\sum_{i=2}^{J+1}\left\{Y_{i}-F_{i}\right\}.$
In the following result, we demonstrate the identifiability of the parameter $\psi_{0}$ in our setting and that $\hat{\psi}^{eq}$ is an unbiased estimator of this parameter.
Theorem 1.
Under Assumptions 1 and 2, the causal parameter $\psi_{0}$ is identified as
$\psi_{0}=\left\{\mathbb{E}[Y_{1}]-\mathbb{E}[F_{1}]\right\}-\frac{1}{J}\sum_{i=2}^{J+1}\left\{\mathbb{E}[Y_{i}]-\mathbb{E}[F_{i}]\right\}.$
See Appendix A.1 for a proof of Theorem 1.
Corollary 1.
Under Assumptions 1 and 2, $\hat{\psi}^{eq1}$ is an unbiased estimator of $\psi_{0}$.
3.2 Logarithmic Equi-Confounding
Suppose outcome variables $F_{i,t}$ are of the same nature as $Y_{i,s}$, but are on a multiplicative scale. Then, clearly, the outcomes in the reference domain can be informative about the potential outcomes in the target domain but the transfer of information can be challenging if the multiplicative scale is not known. For such a scenario, we propose a modification of the linear confounding framework which characterizes the relationship of the variables on a multiplicative scale.
Assumption 3 (Logarithmic equi-confounding).
The potential outcome variables in the target domain and the outcome variables in the reference domain satisfy the following equation.
$\frac{\mathbb{E}[Y_{1}^{(0)}]}{\mathbb{E}[\sum_{i=2}^{J+1}Y_{i}^{(0)}]}=\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\sum_{i=2}^{J+1}F_{i}]}.$ (3)
Similar to Equation (2), in Equation (3), the left hand side is based on the variables in the target domain and the right hand side is based on the variables in the reference domain. Hence, Assumption 3 also specifies the connection between the outcomes in the target domain and the outcomes in the reference domain. We note that, Assumption 3 can be viewed as a modification of Assumption 2, in which the change in mean is replaced by the change in the logarithm of the mean; hence the name logarithmic equi-confounding. A similar observation has been presented in (Wooldridge,, 2023). More specifically, Equation (3) can be written as
$\log\bigg{(}\mathbb{E}[Y_{1}^{(0)}]\bigg{)}-\log\bigg{(}\mathbb{E}[\sum_{i=2}^{J+1}Y_{i}^{(0)}]\bigg{)}=\log\bigg{(}\mathbb{E}[F_{1}]\bigg{)}-\log\bigg{(}\mathbb{E}[\sum_{i=2}^{J+1}F_{i}]\bigg{)}.$
We have the following identification result under the logarithmic equi-confounding assumption.
Theorem 2.
Under Assumptions 1 and 3, the causal parameter $\psi_{0}$ is identified as
$\psi_{0}=\mathbb{E}[Y_{1}]-\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\sum_{i=2}^{J+1}F_{i}]}\mathbb{E}[\sum_{i=2}^{J+1}Y_{i}].$
See Appendix A.2 for a proof of Theorem 2.
Consider the following estimator for causal parameter $\psi_{0}$,
$\hat{\psi}^{eq2}=Y_{1}-\frac{F_{1}}{\sum_{i=2}^{J+1}F_{i}}\sum_{i=2}^{J+1}Y_{i}.$
It can be shown that, under the assumptions of Theorem 2, this estimator is consistent for the parameter $\psi_{0}$. However, we note that in our setting, we only have one observation from the target variable, and hence, such a consistency result is not relevant. Therefore, instead, we aim to bound the bias of the estimator. To this end, we require the following conditions on the variables of the setting.
Assumption 4.
  • For all $i$, $Y_{i}^{(0)}$ is in $[l_{y},L_{y}]$ and $F_{i}$ is in $[l_{f},L_{f}]$, where $0<l_{y},l_{f}$.
  • There exist monotonically decreasing functions $\tau_{1}$ and $\tau_{2}$ with $\tau_{1}(J)\rightarrow 0$ and $\tau_{2}(J)\rightarrow 0$ as $J\rightarrow\infty$, such that
    $\left|\frac{1}{J}\sum_{i=2}^{J+1}cov(F_{1},F_{i})\right|\leq\tau_{1}(J),$
    $\left|\frac{1}{J^{2}}\sum_{i\neq j}cov(F_{i},F_{j})\right|\leq\tau_{2}(J).$
The second part of Assumption 4 is trivially satisfied if the outcome variables in the reference domain are pairwise independent, yet this is not a necessary condition. The example below demonstrates a data generating process in which Assumption 4 is satisfied.
Example 2.
Suppose for all $11\leq i\leq J+$ 1, we have
$Y_{i}^{(0)}=X_{i}+\vartheta_{i}U+\varepsilon_{i},$
$F_{i}=Z_{i}+\theta_{i}U+\epsilon_{i},$
where all variables are bounded, $X_{i}$’s, $Z_{i}$’s, $\varepsilon_{i}$’s, and $\epsilon_{i}$’s are independent variables, $U$, $\varepsilon_{i}$’s, and $\epsilon_{i}$’s are mean-zero, the supports of $X_{i}$’s and $Z_{i}$’s are so that $Y_{i}^{(0)}$’s and $F_{i}$’s are positive and bounded away from zero, and for some functions $\tau_{1}$ and $\tau_{2}$ satisfying the requirements in Assumption 4, $\theta_{i}$’s satisfy $\left|1/J\sum_{i=2}^{J+1}\theta_{i}\right|\leq\tau_{1}(J)$, and $\left|1/J^{2}\sum_{i\neq j}\theta_{i}\theta_{j}\right|\leq\tau_{2}(J)$. It is easy to see that this data generating process satisfies Assumption 4.
Assumption 5.
The variables $F_{1}/\sum_{i=2}^{J+1}F_{i}$ and $\sum_{i=2}^{J+1}Y_{i}$ satisfy
$\left|cov\left(\frac{F_{1}}{\frac{1}{J}\sum_{i=2}^{J+1}F_{i}},\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}\right)\right|\leq C(J),$
for some monotonically decreasing function for which $C(J)$ converges to zero as $J\rightarrow\infty$.
Regarding Assumption 5, note that this assumption does not require the outcome variables in the target and the reference domains to be (approximately) uncorrelated, but it requires that for $1/J\sum_{i=2}^{J+1}Y_{i}$ and the ratio quantity $F_{1}/(1/J\sum_{i=2}^{J+1}F_{i})$, which is a much weaker requirement, as the ration can remain approximately unchanged even when variables are changing.
We have the following result regarding bounding the bias of our proposed estimator.
Theorem 3.
Under Assumptions 1, 3, 4 and 5, the absolute bias of the estimator $\hat{\psi}^{eq2}$ can be bounded as
$\left|\psi_{0}-\mathbb{E}[\hat{\psi}^{eq2}]\right|\leq L_{y}\left[\frac{l_{f}\tau_{1}(J)+L_{f}\tau_{2}(J)+L_{f}\tau_{2}(J)+L_{f}L_{y}^{2}/J}{l_{f}^{3}}\right]+C(J).$
See Appendix A.3 for a proof of Theorem 3. This result demonstrates that under the assumptions required on the data generating process, the bound on the bias of our proposed estimator vanishes as the number of units that we compare with the target unit increases.
As it can be seen from the results presented in this section, the equi-confounding data fusion framework allows for simple identification and estimation processes. However, the assumption of this framework may rule out certain data generating processes with time-unit interaction terms. In the next section, we present an alternative data fusion approach, which allows for more flexible data generating processes.
4 Synthetic Control Data Fusion
The equi-confounding assumptions introduced in the previous section require a relation between the average of the outcomes for the control units and the outcomes of the target unit in the reference and the target domains. In this averaging, all units are weighted equally, with weights $1/J$. In many applications, certain control units are more similar to the target unit and hence, may contain more information regarding the counterfactual outcomes of the target unit. Therefore, a natural extension of the ideas in the previous section is to consider different weights for different control units. This is in fact the main idea of the synthetic control method (Abadie and Gardeazabal,, 2003; Abadie et al.,, 2010, 2015). Our goal is to formalize this idea in our data fusion setting.
We consider weight $w_{i}$ for the $i$-th control unit, $12\leq i\leq J+$ 1, and we define the $J$-dimensional weight vector $w=[w_{2}~{}\cdots~{}w_{J+1}]^{\top}$, such that the weights must satisfy
$w_{i}\geq 0,~{}~{}~{}~{}~{}i=2,\dots,J+1,~{}~{}~{}\text{and}~{}~{}~{}\sum_{i=2}^{J+1}w_{i}=1.$
Our goal is to choose weights such that the following restrictions are (approximately) satisfied.
  • The weighted sum of the outcomes of the control units in the reference domain at time $t$ matches that of the target unit, for $1\leq t\leq T$. That is,
    $F_{1t}=\sum_{i=2}^{J+1}w_{i}F_{it},~{}~{}~{}~{}~{}~{}1\leq t\leq T.$ (4)
  • The weighted sum of the observed covariates of the control units in the reference domain matches that of the target unit. That is,
    $Z_{1}=\sum_{i=2}^{J+1}w_{i}Z_{i}.$ (5)
  • The weighted sum of the observed covariates of the control units in the target domain matches that of the target unit. That is,
    $X_{1}=\sum_{i=2}^{J+1}w_{i}X_{i}.$ (6)
The intuition of imposing the above restrictions is as follows. Suppose the above matching requirements are satisfied, and also, the outcomes of the units in the reference and target domains are generated by mechanisms that are mostly only different in the way they are affected by the observed covariates. In this case, if we are matching on the said observed covariates and the outcomes on a long period of time in the reference domain, then we should have been also matching on any unobserved covariate affecting the outcomes in both domains. For instance, for the application of vaccination rate with the choice of the Hispanic and Black sub-populations as the target and reference domains, respectively, peoples’ overall attitude towards vaccines, socioeconomic status, and access to vaccines are unobserved covariates that may similarly affect COVID-19 vaccination rates across the two sub-populations. Therefore, if we match on vaccination rate from one subpopulation for an extended time period, as well as all the observed covariates concerning for COVID-19 vaccination, then there is statistical confidence that we are matching on the unobserved covariates as well.
Based on the above argument, a weighted sum of the potential outcomes with weight vector $w$ that (approximately) satisfies restrictions (4)-(6) will match the potential outcome of the target unit with a good approximation. That is,
$Y_{1}^{(0)}\approx\sum_{i=2}^{J+1}w_{i}Y_{i}^{(0)}=\sum_{i=2}^{J+1}w_{i}Y_{i},$
where the last equality follows from the consistency assumption. We refer to the weighted sum on the right hand side as the transferred synthetic control. Equippied with this quantity, we propose the following estimator for our causal parameter of interest:
$\hat{\psi}^{sc}=Y_{1}-\sum_{i=2}^{J+1}w_{i}Y_{i}.$ (7)
We note that in restrictions (4)-(6), we will usually have more equations than unknowns, and hence, these restrictions cannot be exactly satisfied. Therefore, we instead, minimize the miss-match in each restriction. Formally, for $i=1,\dots,J+1$, let $\bar{F}_{i}:=[F_{i,1}~{}\cdots~{}F_{i,T}]^{\top}$, and $K_{i}:=[\bar{F}_{i}^{\top}~{}Z_{i}^{\top}~{}X_{i}^{\top}]^{\top}$. A naive way to obtain weight vector $w$ is as follows:
$\begin{array}[]{ll}w=\arg\min_{w}&\left\|K_{1}-\sum_{i=2}^{J+1}w_{i}K_{i}\right\|_{2}^{2},\\ {}\\ \text{~{}~{}~{}~{}~{}~{}~{}~{}~{}s.t.}&\sum_{i=2}^{J+1}w_{i}=1,\\ &w_{i}\geq 0,\quad i=2,\ldots,J+1.\end{array}$ (8)
This is in fact akin to the methodology used in the original synthetic control method. However, unfortunately, this basic approach does not suffice for our data fusion goal. An important concern with the approach above is that if we have a larger $T$ than $d_{t}$, the chosen weight vector $w$ may be mostly influenced by $\bar{F}_{i}$, resulting in a good matching in the reference domain but poor tracking of the $X_{i}$ in the target domain, and hence, poor matching on the potential outcomes in the target domain. Note that this issue does not occur on the standard synthetic control setting because in that setting, matching on the outcomes help better matching on the observed covariates and vice versa. But in our data fusion setting, matching on observed variables in one domain does not necessarily improve that in the other domain. In order to quantify the performances of the matchings, for a given $w$, we define the normalized squared errors (NSEs) of the predictors of $\bar{F}_{1}$, $Z_{1}$, and $X_{1}$, as
$NSE(O_{1},w):=\frac{1}{|O_{1}|}\times\left\|{O_{1}-\sum_{i=2}^{J+1}w_{i}O_{i}}\right\|_{2}^{2},$
where $O_{i}$ could be $\bar{F}_{i}$, $Z_{i}$, or $X_{i}$.
To resolve the issue above, we add an extra constraint to the optimization problem that forces the matching on the observed covariates to be close to the best possible matching. We further incorporate a budgeting vector $b$ into our objective function to achieve more flexibility in the choice of the weights. Formally, let $b=[b_{F}~{}b_{Z}~{}b_{X}]^{\top}$ be a vector with positive elements that sum up to one, that determines the budgets that we will spend on minimizing the NSEs for $\bar{F}_{1}$, $Z_{1}$, and $X_{1}$. This vector can be simply set to $b=[1/3~{}1/3~{}1/3]^{\top}$. Alternatively, we can choose to prioritize minimizing the NSE of $\bar{F}_{1}$, $Z_{1}$, or $X_{1}$ in case we have knowledge that it is more complex than the other two by increasing its corresponding element in $b$. In general, this can be done in a data driven fashion using estimated NSEs. We will next propose an algorithm that chooses the budget vector $b$ in a data-driven approach, and uses it to find an optimal weight vector $w^{*}$.
Our method for choosing the optimal budget vector $b^{*}$ and weights $w^{*}$ is outlined in Algorithm 1. This algorithm takes the observed data as well as two hyper-parameters $\eta_{Z}$ and $\eta_{X}$ as the input. We first calculate baseline weights $w^{*}_{Z}$ and $w^{*}_{X}$ separately by minimizing $NSE(Z_{1},w)$ and $NSE(X_{1},w)$, respectively. These baseline weights serve as reference points for constraining the optimization process, as they lead the best possible NSEs that one can get for $X_{1}$ and $Z_{1}$. Next, we define $B$ as the set of all possible vectors $b$ satisfying $b_{F},b_{Z},b_{X}\geq 0$ and $b_{F}+b_{Z}+b_{X}=1$. For each vector $b\in B$, the algorithm finds $w^{*}(b)$ by solving optimization problem 9. This optimization problem aims to find $w$ which minimizes the weighted sum of NSEs. Importantly, we require the solution to be such that the NSEs for $Z_{1}$ and $X_{1}$ are not far away from the best possible NSEs that one can get for these two variables (using baseline weights $w^{*}_{Z}$ and $w^{*}_{X}$). This ensures that the matching for the covariates will not be compromised, and hence resolves the concern mentioned earlier. Finally, the algorithm picks $b^{*}$ by minimizing $NSE(\bar{F}_{1},w^{*}(b))$, and returns $w^{*}=W^{*}(b^{*})$. Therefore, among all the choices with acceptable match on the covariates, we pick the one with the best match on the outcomes in the reference domain. This choice ensures a good match on unobserved components involved in the generating mechanism of the outcomes.
Input: $(\bar{F}_{i},Z_{i},X_{i})_{i=1}^{J+1}$, and hyper-parameters $\eta_{Z}$, $\eta_{X}$.
Define: $B=\{b=\{b_{F},b_{Z},b_{X}\}:b_{F},b_{Z},b_{X}\geq 0\text{ and }b_{F}+b_{Z}+b_{X}=1\}$.
Calculate:
$w^{*}_{Z}=\arg\min_{w}NSE(Z_{1},w),$ s.t. $\sum_{i}w_{i}=1,\quad 0\leq w_{i}$;
$w^{*}_{X}=\arg\min_{w}NSE(X_{1}.w)$ s.t. $\sum_{i}w_{i}=1,\quad 0\leq w_{i}$.
for $b\in B$ do
       Solve:
$w^{*}(b)$ $=\arg\min_{w}b_{F}\cdot NSE(\bar{F}_{1},w)+b_{Z}\cdot NSE(Z_{1},w)+b_{X}\cdot NSE(X_{1},w)$ (9)
s.t. $\quad\frac{1+NSE(Z_{1},w)}{1+NSE(Z_{1},w^{*}_{Z})}\leq 1+\eta_{Z},\quad\frac{1+NSE(X_{1},w)}{1+NSE(X_{1},w^{*}_{X})}\leq 1+\eta_{X},$
$\quad\sum_{i}w_{i}=1,\quad 0\leq w_{i}.$
end for
Set $b^{*}=\arg\min_{b\in B}NSE(\bar{F}_{1},w^{*}(b))$.
Return $w^{*}=w^{*}(b^{*})$.
Algorithm 1 Synthetic control data fusion algorithm
Remark 1.
The objective function in the optimization problem (9) can be written in a more concise way as follows. Define $\lambda_{F}(b_{F})=\sqrt{b_{F}/T}$, $\lambda_{Z}(b_{Z})=\sqrt{b_{Z}/d_{r}}$, and $\lambda_{X}(b_{X})=\sqrt{b_{X}/d_{t}}$. Using this notation, define $K_{i}^{\lambda}:=[\lambda_{F}(b_{F}^{*})\bar{F}_{i}^{\top}~{}\lambda_{Z}(b_{Z}^{*})Z_{i}^{\top}~{}\lambda_{X}(b_{X}^{*})X_{i}^{\top}]^{\top}$. Then the objective function can be written as $\|K_{1}^{\lambda}-\sum_{i=2}^{J+1}w_{i}K_{i}^{\lambda}\|^{2}_{2}$. This form is closer to the form of the objective function in (8). However, note that as mentioned earlier, we have extra constrains in optimization problem (9) compared to (8).
Remark 2.
In the approach above, in line with the original synthetic control method, we require the weights $w$ to be non-negative and that they sum up to one. The purpose of this restriction is interpretability and avoiding over-fitting. Yet, similar to the standard synthetic control method, we can also prevent over-fitting by regularization (Doudchenko and Imbens,, 2016).
As mentioned earlier, our proposed method for approximating the counterfactual outcome of the target unit in the target domain is expected to have a reasonable performance if the outcomes of the units in the reference and target domains are generated by mechanisms that are mostly only different in the way they are affected by the observed covariates. We formalize this requirement by considering the following structural equations for the outcome variables in the reference and target domains.
Assumption 6.
The outcome variables in the reference and target domains are generated from the following factor models. For all $11\leq i\leq J+$ 1, $1\leq s\leq S$, and $1\leq t\leq T$,
$Y_{i,s}=\varrho_{s}+\varphi_{s}^{\top}X_{i}+\vartheta_{s}^{\top}\mu_{i}+\alpha_{i,s}A_{i,s}+\varepsilon_{i,s},$ (10)
$F_{i,t}=\rho_{t}+\phi_{t}^{\top}Z_{i}+\theta_{t}^{\top}\mu_{i}+\epsilon_{i,t}.$
$\varrho_{s}$, $\rho_{t}$, $\varphi_{t}$, $\phi_{t}$, $\vartheta_{t}$, $\theta_{t}$, and $\mu_{i}$ are conformable fixed but unknown real-valued factors. $\alpha_{i,s}$ is the causal effect on unit $i$ in time period $s$. $\mu_{i}$ is an $d_{u}$-dimensional latent factor that is specific to unit $i$ and affects its outcome in both domains. We assume that for all $s$, $t$, and $l$, $|\theta_{t,l}|\leq\overline{\theta}$, and $|\vartheta_{s,l}|\leq\overline{\vartheta}$. Moreover, the eigenvalues of the matrix $\sum_{l=1}^{T}\theta_{l}{\theta_{l}}^{\top}$ are lower bounded by $T\underline{\xi}>0$. $\epsilon_{i,t}$ and $\varepsilon_{i,s}$ are unobserved transitory shocks at the unit level with zero mean. $\epsilon_{i,t}$’s are i.i.d. sub-Gaussian random variables with variance proxy $\bar{\sigma}^{2}$.
In Assumption 6, we require that the unobserved unit-specific factors ($\mu_{i}$’s) be the same in the two factor models. For example, in the vaccination rate application mentioned earlier, $\mu_{i}$ is the overall attitude and belief regarding vaccination for the $i$-th individual. Note that under the factor model (10), the causal parameter of interest can be written as
$\psi_{0}=\frac{1}{S}\sum_{s=1}^{S}\alpha_{1,s}.$
We have the following result regarding bounding the bias of our proposed estimator.
Theorem 4.
Suppose Conditions (4)-(6) hold. Under the Assumptions 1 and 6, we have
$\left|\mathbb{E}[\hat{\psi}^{sc}]-\psi_{0}\right|\leq\sqrt{2}\bar{\vartheta}\bar{\theta}d_{u}\bar{\sigma}\underline{\xi}^{-1}\sqrt{\frac{\log(2J)}{T}}.$
See Appendix A.4 for a proof of Theorem 4.
Theorem 4 demonstrates that the bound on the bias of our proposed estimator vanishes as the number of time periods in the reference domain increases. We also see that the bound becomes wider as the number of donor units increases. This is due to the fact that with more donor units, there is a higher chance of overfitting, i.e., matching the units by chance. A main ingredient in the proof is to use a maximal inequality to bound a linear combination of weighted error terms. This step was also required in the work of Abadie et al., (2010) in the standard synthetic control setting. However, compared to that work, we took a different approach and used a methodology based on the assumption of sub-Gaussianity of the noise terms (See Assumption 6). This will lead to a more clear and interpretable derivation with no need for assumption on moments of certain variables used in (Abadie et al.,, 2010). Besides this difference, we note that we also have to deal with two sets of factor loadings for the latent factors $\mu_{i}$ corresponding to the two domains. Hence, we require assumptions for bounding both sets. It can be seen in Theorem 4 that the upper bound on both factor loadings appear in the bound on the bias.
In the next section, we evaluate the performance of our proposed estimator in a simulation study, and compare it with the estimators introduced in Section 3.
5 Simulation Studies
In this section, we present the results of the evaluation of the performance of our proposed estimation methods. We considered a setting with 31 units ($J=30$) where we observe $d_{r}=3$ covariates in the reference domain and $d_{t}=3$ covariates in the target domain. Moreover, we considered 3 latent confounders (i.e., $d_{u}=3$) in the system which affect the outcomes of the units in both domains. We observe $T\in\{10,20,...,100\}$ time periods in the reference domain and $S=5$ time periods in the target domain. We generated the data from the factor model in Display (10).
For all $i$, $s$, and $t$, entries of $X_{i}$, $Z_{i}$, and $\mu_{i}$ were generated i.i.d. from Unif$(0,1)$, entries of $\phi_{t}$, $\theta_{t}$, $\varphi_{s}$, and $\vartheta_{s}$ were generated i.i.d. from Unif$(0,10)$. $\rho_{t}$ was generated i.i.d. from Unif$(0,20)$ and sorted in the increasing order. The same process was used to generate $\varrho_{s}$ but with Unif$(0,10)$. $\epsilon_{i,t}$ and $\varepsilon_{i,s}$ were generated i.i.d. from Gaussian distribution $\mathcal{N}(0,2)$ and $\mathcal{N}(0,0.5)$, respectively. Finally, the parameters $\alpha_{1,s}$ corresponding to the causal effect on the target unit were generated i.i.d. from Unif$(2,5)$ and sorted in the increasing order. Fixing the factors in the data generating process (DGP) from the above distributions, in order to evaluate our methods, we generated $M=300$ datasets from the DGP. In order to see the effect of increasing $T$ from $T=T_{1}$ to a larger value $T=T_{2}$, the first $T_{1}$ factors in the latter are the same as the $T_{1}$ factors in the former. For an estimator $\hat{\psi}_{T}$ operating over $T$ time periods in the reference domain, we estimated the bias using $M$ datasets as $1/M\sum_{m=1}^{M}\hat{\psi}_{T}(D_{m})-\psi_{0}$, where $D_{m}$ is the $m$-th dataset.
(a) Evaluation of linear equi-confounding assumption on one DGP.
(b) Evaluation of logarithmic equi-confounding assumption on one DGP.
Figure 1: Visual demonstration of the equi-confounding assumptions.
For the linear equi-confounding approach, let $F_{0}=\frac{1}{J}\sum_{i=2}^{J+1}{F_{i}}$ and $Y_{0}=\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}$. Note that Assumption 2 does not necessarily hold for the factor model described in Equation 10. For instance, Figure 1(a) demonstrates one realization of the DGP ($T=20$) in which the line connecting $\mathbb{E}[F_{1}]$ to $\mathbb{E}[Y_{1}^{(0)}]$ is not parallel to the line connecting $\mathbb{E}[F_{0}]$ to $\mathbb{E}[Y_{0}]$. Hence, the estimator $\hat{\psi}^{eq1}$ is expected to be biased. However, over extensive simulations (several realizations of the DGP), we observed that with our choice of distributions and parameters, the bias normalized by ground truth casual effect in most of the cases was negligible. Similar concern holds for the logarithmic equi-confounding approach, that is, Assumption 3 does not necessarily hold. Specifically, the lines connecting $\log(\mathbb{E}[F_{1}])$ to $\log(\mathbb{E}[Y_{1}^{(0)}])$ and $\log(\mathbb{E}[F_{0}])$ to $\log(\mathbb{E}[Y_{0}])$ are not parallel (see Figure 1(b)). However, comparing Figures 1(a) and 1(b), in the corresponding DGP we see smaller violation of the logarithmic equi-confounding assumption.
(a) Units in the reference domain and the synthetic control target unit.
(b) Units in the target domain, the synthetic control target unit, and the counterfactual target unit.
Figure 2: Factual, counterfactual, and estimated units in the synthetic control approach in the reference and target domains.
Moving to the synthetic control method, we notice that the optimization problem in Equation 9 takes the form of a quadratic objective function with quadratic constraints. To solve it, we use the gurobi package (Gurobi Optimization, LLC,, 2024) in the software R. One key point of our methodology is the tradeoff between the matching qualities of the reference and the target domains. This balance is governed by the flexibility in terms of the matching quality we allow by the choice of the parameters $\eta_{Z}$ and $\eta_{X}$, which control the relative importance of matching in the reference and target domains, respectively. A larger $\eta_{Z}$ allows for better matching in the target domain at the expense of the reference domain, whereas a larger $\eta_{X}$ allows for a better alignment in the reference domain but may reduce the quality of the matching in the target domain. This tradeoff can be adjusted by the user based on domain-specific considerations in terms of the association between the covariates and the outcome variables, and estimation priorities. In our implementation, we set both $\eta_{Z}$ and $\eta_{X}$ to 0.1. However, we also performed sensitivity analysis to investigate the sensitivity of our estimated causal effect to the choice of $\eta_{Z}$ and $\eta_{X}$. The result of sensitivity analysis is presented in Figure 5(a). As it can be seen in figure, deviations from the chosen value of $0.1$ either increasing or decreasing $\eta_{Z}$ and $\eta_{X}$-exhibit minimal impact on the causal effects.
The same DGP described above was used for synthetic control method. We drew a single dataset with $T=20$. In Figure 2(a), the black solid line shows the trajectory of the observed target unit and the black dashed line shows the trajectory of $\sum_{i=2}^{J+1}w^{*}_{i}F_{it}$ for $t\in\{1,\dots,T\}$. The grey solid lines represent the observed control units for $i\in\{2,\dots,J+1\}$. We observe that the synthetic control target unit closely tracks the observed target unit in the reference domain. In Figure 2(b), the black and grey solid lines represent the paths of the observed units for $i\in\{1,\dots,J+1\}$ and $s\in\{1,\dots,S\}$, and the solid red line represents the counterfactual target unit in the target domain. The dashed line shows the synthetic control target unit $\sum_{i=2}^{J+1}w^{*}_{i}Y_{i,s}$ using the same weights $w^{*}$ mentioned above. Figure 2(b) demonstrates that the synthetic control data fusion method is able to provide a good fit for $Y_{1,s}^{(0)}$ using the weights $w^{*}$ learned from the reference domain.
To demonstrate the matching performance from our proposed method, we calculated the NSEs for latent confounders, covariates, and outcomes in the reference domain. Importantly, the NSE for $\bar{F}_{1}$ starts from a lower value and then converges to a slightly higher value. This is due to the fact that with shorter time period, we will have over-fitting and the algorithm matches on the errors, but with increased length of the time period, the over-fitting is resolved. Note that, the NSE does not converge to zero, because there will be noise for the outcome variable in every time period and hence one should not expect to have an NSE smaller than the variance of the noise. One the other hand, the NSE decreases for $Z_{1}$ because as the length of the time period increases more matching on $\bar{F}_{1}$ helps better matching on $Z_{1}$. This indicates that the estimated matches become more accurate with more data points. Note that we do not directly target for a good matching for $\mu_{1}$ in the optimization problem. Yet the optimized weights $w^{*}$ yield a good match between $\mu_{1}$ and $\sum_{i=2}^{J+1}w^{*}_{i}\mu_{i}$, as seen in Figure 3(b).
(a) NSEs of $\mu_{i}$, $Z_{i}$, $X_{i}$, and $\bar{F}_{i}$.
(b) NSEs of $\mu_{i}$, $Z_{i}$, and $X_{i}$.
(c) Differences of $\hat{\psi}^{eq1}$, $\hat{\psi}^{eq2}$, $\hat{\psi}^{sc}$ from $\psi_{0}$.
Figure 3: Bias and NSEs in synthetic control and equi-confounding methods.
Figure 3(c) provides a comparative visualization of the biases from the synthetic control data fusion and equi-confounding data fusion methods. We present the biases calculated by averaging over 300 datasets for each value of $T$, along with the interquartile range. Overall, we observed a superior performance from the synthetic control data fusion method. The bias from the synthetic control data fusion method decreases as the value of $T$ increases, eventually approaching zero. This reaffirms the result in Theorem 4 that the bias approaches zero as the number of time periods increases. The logarithmic equi-confounding method also presents a decreasing bias trend. However, there is no decreasing pattern for bias from the linear equi-confounding method.
(a) Placebo test for reference domain.
(b) Placebo test for target domain.
Figure 4: Placebo test.
(a) Sensitivity analysis from the simulated data.
(b) Sensitivity analysis from the real data implementation.
Figure 5: Sensitivity analysis from simulation and application.
5.1 Placebo Test
To evaluate the ability of the synthetic control method in predicting the counterfactual outcome, we implemented placebo test, the inferential technique originated from (Abadie and Gardeazabal,, 2003). We implemented the placebo method by applying the synthetic control method to each unit in the donor pool. We generated synthetic controls for each donor using the same procedure as for the target unit and shifting the target unit to the donor pool. We then compared the distribution of these placebo effects to the effect estimated for the target unit. By comparing the distribution of estimated placebo effects to the effect estimated for the target unit, we can determine whether the observed effect is significantly larger than what would be expected by chance, thereby strengthening our inference about the causal impact of the intervention.
The results of the placebo test are shown in Figure 4. The grey lines show the difference between each of the 30 units in the donor pool and their synthetic controls. While the black line represents that for the target unit. Figure 4(a) presents that the placebo effects are centered around zero, indicating negligible bias in the estimation process in the reference domain. From Figure 4(b), it can be observed that the estimated difference for the target unit and its synthetic control is significantly higher compared to that for the donors, demonstrating statistical significance of the estimated positive causal effect of the treatment.
Synthetic Covariates Chelsea Chelsea Average of 19 controls Median income - Hispanic 0.522 0.388 0.367 Median income - Black 0.206 0.371 0.388 Proportion - Hispanic 0.784 0.479 0.294 Proportion - Black 0.362 0.350 0.310 Median age - Hispanic 0.73 0.526 0.396 Median age - Black 0.418 0.414 0.406 Proportion aged 65+ - Hispanic 0.637 0.348 0.265 Proportion aged 65+ - Black 0.230 0.309 0.235
Table 1: Covariate means for Chelsea, synthetic Chelsea, and average of control units.
6 Application to COVID-19 Vaccination Rate Analysis
During the COVID-19 pandemic, community organizations played a pivotal role in promoting health equity through localized efforts. In our empirical study, we evaluated the effect of the pro-vaccination initiatives of La Colaborativa, a major Hispanic community-based organization in Chelsea, Massachusetts, on the COVID-19 vaccination rate of the Hispanic subpopulation of that city. Chelsea was among the 20 municipalities in Massachusetts most severely impacted by COVID-19. This disproportionate impact was evident in both case rates and social determinants of health, particularly among Hispanic (which comprises $60\%$ of Chelsea’s population) and Black sub-populations.
Despite the urgent need for vaccination, access barriers in the early rollout disproportionately affected non-English-speaking and low-income populations. When access first became available in Massachusets, eligibility alone did not guarantee access–many residents faced challenges such as language barriers, lack of reliable transportation, and difficulties securing online appointments. These disparities created an urgent need for targeted community-based interventions to ensure equitable vaccine distribution. To address this challenge, La Colaborativa launched a series of pro-vaccination efforts in March 2021, in alignment with the Massachusetts Vaccine Equity Initiative. These initiatives included door-to-door outreach, mobile vaccination clinics, and strategic partnerships with the East Boston Neighborhood Health Center to increase accessibility for Chelsea’s Hispanic population.
Evaluating the impact of these interventions poses a significant challenge due to the timing of the available data. The Massachusetts Public Health Department began recording COVID-19 vaccination rates in March 2021-coinciding with the start of La Colaborativa’s efforts. Hence, no pre-intervention data were collected. This lack of a pre-intervention period prevents the use of traditional causal inference approaches, which typically rely on pre-treatment trends to establish counterfactual comparisons. To address this limitation, we applied our proposed methodology to estimate the causal effects of La Colaborative’s intervention.
We considered the Hispanic sub-population in 20 cities in Massachusetts, as our target domain with monthly COVID-19 vaccination rate from March 2021 to July 2022 as the outcome variable, and with covariates median income, proportion of the subpopulation, median age, and proportion of residents aged 65 or older of the Hispanic sub-population. The city of Chelsea, which is our target unit, is among these 20 cities. The rest of the 20 cities were chosen based on similarities in their covariates to Chelsea. Importantly, cities served by similar community organizations to La Colaborativa were excluded from the donor pool to ensure a valid comparison group. We considered the Black sub-population in the same 20 cities as our reference domain. The reasoning behind this choice for the reference domain was that both Black and Hispanic populations face systemic barriers to healthcare access, vaccine hesitancy, and similar structural inequalities. These shared factors make Black vaccination trends a valuable reference for estimating causal effects in the Hispanic population.
Unit Weight Unit Weight
Bellingham 0.000 Marlborough 0.000
Chicopee 0.000 Methuen 0.000
Everett 0.228 Milford 0.000
Fitchburg 0.000 New Bedford 0.000
Haverhill 0.000 Peabody 0.000
Holyoke 0.310 Revere 0.330
Lawrence 0.037 Southbridge 0.000
Leominster 0.000 Springfield 0.094
Lynn 0.000 Waltham 0.000
Worcester 0.000
Table 2: Unit weights in the synthetic Chelsea.
(a) COVID-19 vaccination rates for Black sub-population (reference domain).
(b) COVID-19 vaccination rates for Hispanic sub-population (target domain).
Figure 6: COVID-19 vaccination rates for the observed units and synthetic Chelsea in reference and target domains.
(a) Placebo test for COVID-19 vaccination rates for Black sub-population (reference domain).
(b) Placebo test for COVID-19 vaccination rates for Hispanic sub-population (target domain).
Figure 7: Placebo test for COVID-19 vaccination rate analysis.
Results:
We used the value 0.1 for both $\eta_{Z}$ and $\eta_{X}$. However, as in the simulation study, we performed sensitivity analysis to investigate the sensitivity of our estimated causal effect to the choice of $\eta_{Z}$ and $\eta_{X}$, and we observed small impact of deviating these hyper-parameters (see Figure 5(b)). Table 1 presents the covariates used in the estimation process. As it can be seen, the method successfully generated the synthetic Chelsea with covariates close to those of the real Chelsea. Table 2 displays the weights assigned to each control city in constructing synthetic Chelsea. The results indicate that the best matching for Chelsea is achieved by combining Revere, Holyoke, Everett, Lawrence, and Springfield. All other units in the donor pool were assigned zero weights.
Figure 6 presents the COVID-19 vaccination rates for the observed units and synthetic Chelsea in reference and target domains. In Figure 6(a), synthetic Chelsea closely follows the actual Chelsea in the reference domain, demonstrating that our algorithms has succeeded to obtain a good match. This, along with the high degree of balance in observed covariates, supports the validity of synthetic Chelsea as a counterfactual for actual Chelsea in the absence of the intervention. In Figure 6(b), the observed Chelsea Hispanic COVID-19 vaccination rate and its synthetic version begin to diverge after the intervention starts, with the gap widening over time. Specifically, it can be seen that Chelsea has a steeper rise in the earlier months. The divergence between the black solid line (actual Chelsea) and the black dashed line (synthetic Chelsea) suggests that La Colaborativa’s efforts contributed to a greater increase in Hispanic vaccination rates than would have been expected without the intervention. We have also provided the results for placebo test in Figure 7(a), which confirms the statistical significance of our findings regarding the causal effect.
We now present the estimated causal effect of the community organization intervention based on our proposed methodologies. Over the 16-month period from March 2021 to July 2022, the COVID-19 vaccination rates for Hispanic sub-population was estimated to be increased by $13.6\%$ using the linear equi-confounding method, $13.2\%$ using the logarithmic equi-confounding method, and $10.1\%$ using the synthetic control data fusion method. All three methods indicate a similar large positive causal effect, which affirms the effectiveness of the community organization intervention.
7 Conclusion
We introduced two data fusion-based methods for identifying and estimating causal effects in panel data settings where no pre-intervention data are available. We introduced the equi-confounding data fusion and synthetic control data fusion methods, through which we addressed the challenges posed by sudden events that leave researchers without the typical pre-intervention period required by conventional causal inference panel data methods. We showed the efficacy of these methods by obtaining bounds on the biases, and provided conditions under which the bounds converge to zero. We showed that the synthetic control data fusion method is more flexible as it requires generally weaker assumptions. On the other hand, the equi-confounding methods are simpler to implement and to interpret. Our simulation results confirm our theoretical findings and demonstrate the fast decreasing rate of the biases. Our proposed causal panel data fusion methods have high potentials to be used in situations where sudden shocks preclude the collection of pre-intervention data. As an example of such applications, we applied our proposed methods to estimate the causal effect of a community organization vaccination intervention in Chelsea, Massachusetts. Our results demonstrate that over the 16-month period from March 2021 to July 2022, the Hispanic COVID-19 vaccination rate in Chelsea increased by $13.6\%$ using the linear equi-confounding method, $13.2\%$ using the logarithmic equi-confounding method, and $10.1\%$ using the synthetic control data fusion method. All three methods indicate a similar large positive causal effect, which affirms the effectiveness of the community organization intervention.
Acknowledgements
We would like to express our gratitude to the Massachusetts Department of Public Health (DPH). We extend special thanks to Boudu Bingay, Elizabeth T. Russo, and Joshua Norville, for their support throughout this project.
Appendix A Proofs
A.1 Proof of Theorem 1
Due to the consistency assumption, $\psi_{0}$ can be written as
$\psi_{0}=\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}[Y_{1,s}-Y_{1,s}^{(0)}]=\mathbb{E}[Y_{1}]-\mathbb{E}[Y_{1}^{(0)}].$
By Assumptions 2,
$\mathbb{E}[Y_{1}^{(0)}]$ $=\frac{1}{J}\sum_{i=2}^{J+1}\mathbb{E}[Y_{i}^{(0)}]+\mathbb{E}[F_{1}]-\frac{1}{J}\sum_{i=2}^{J+1}\mathbb{E}[F_{i}].$
Therefore,
$\psi_{0}$ $=\mathbb{E}[Y_{1}]-\mathbb{E}[Y_{1}^{(0)}]$
$=\mathbb{E}[Y_{1}]-\frac{1}{J}\sum_{i=2}^{J+1}\mathbb{E}[Y_{i}^{(0)}]-\mathbb{E}[F_{1}]+\frac{1}{J}\sum_{i=2}^{J+1}\mathbb{E}[F_{i}].$
Finally, applying the consistency assumption one more time concludes the desired result.
A.2 Proof of Theorem 2
Due to the consistency assumption, $\psi_{0}$ can be written as
$\psi_{0}=\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}[Y_{1,s}-Y_{1,s}^{(0)}]=\mathbb{E}[Y_{1}]-\mathbb{E}[Y_{1}^{(0)}].$
By Assumption 3,
$\mathbb{E}[Y_{1}^{(0)}]$ $=\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\sum_{i=2}^{J+1}F_{i}]}\mathbb{E}[\sum_{i=2}^{J+1}Y_{i}^{(0)}].$
Therefore, applying the consistency assumption one more time, we obtain
$\psi_{0}$ $=\mathbb{E}[Y_{1}]-\mathbb{E}[Y_{1}^{(0)}]$
$=\mathbb{E}[Y_{1}]-\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\sum_{i=2}^{J+1}F_{i}]}\mathbb{E}[\sum_{i=2}^{J+1}Y_{i}].$
A.3 Proof of Theorem 3
From Theorem 2, the parameter of interest $\psi_{0}$ is identified as
$\psi_{0}=\mathbb{E}[Y_{1}]-\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\sum_{i=2}^{J+1}F_{i}]}\mathbb{E}\left[\sum_{i=2}^{J+1}Y_{i}\right].$
For the estimator $\hat{\psi}^{eq2}$, we have
$\mathbb{E}[\hat{\psi}^{eq2}]$ $=\mathbb{E}[Y_{1}]-\mathbb{E}\left[\frac{F_{1}}{\sum_{i=2}^{J+1}F_{i}}\sum_{i=2}^{J+1}Y_{i}\right].$
Hence, using Assumption 5, the absolute bias of $\hat{\psi}^{eq2}$ will be bounded as
$\left|\psi_{0}-\mathbb{E}[\hat{\psi}^{eq2}]\right|$
$=\left|\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\sum_{i=2}^{J+1}F_{i}]}\mathbb{E}\left[\sum_{i=2}^{J+1}Y_{i}\right]-\mathbb{E}\left[\frac{F_{1}}{\sum_{i=2}^{J+1}F_{i}}\right]\mathbb{E}\left[\sum_{i=2}^{J+1}Y_{i}\right]+cov\left(\frac{F_{1}}{\frac{1}{J}\sum_{i=2}^{J+1}F_{i}},\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}\right)\right|$
$\leq\left|\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}]}\mathbb{E}\left[\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}\right]-\mathbb{E}\left[\frac{F_{1}}{\frac{1}{J}\sum_{i=2}^{J+1}F_{i}}\right]\mathbb{E}\left[\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}\right]\right|+C(J)$
$=\left|\mathbb{E}\left[\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}\right]\left\{\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}]}-\mathbb{E}\left[\frac{F_{1}}{\frac{1}{J}\sum_{i=2}^{J+1}F_{i}}\right]\right\}\right|+C(J)$
$\leq L_{y}\left|\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}]}-\mathbb{E}\left[\frac{F_{1}}{\frac{1}{J}\sum_{i=2}^{J+1}F_{i}}\right]\right|+C(J),$
where the last inequality follows from Assumption 4.
By a Taylor expansion around $\mathbb{E}[F_{1}]/\mathbb{E}[1/J\sum_{i=2}^{J+1}F_{i}]$, we have
$\mathbb{E}\left[\frac{F_{1}}{\frac{1}{J}\sum_{i=2}^{J+1}F_{i}}\right]=~{}$ $\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}]}$
$+\frac{1}{\mathbb{E}[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}]}\mathbb{E}[F_{1}-\mathbb{E}[F_{1}]]$
$-\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}]^{2}}\mathbb{E}\left[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}-\mathbb{E}\left[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}\right]\right]$
$+\frac{\mu_{1}}{\mu_{0}^{3}}\mathbb{E}\left[\left(\frac{1}{J}\sum_{i=2}^{J+1}F_{i}-\mathbb{E}\left[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}\right]\right)^{2}\right]$
$-\frac{1}{\mu_{0}^{2}}\mathbb{E}\left[\Bigg{(}F_{1}-\mathbb{E}[F_{1}]\Bigg{)}\left(\frac{1}{J}\sum_{i=2}^{J+1}F_{i}-\mathbb{E}\left[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}\right]\right)\right],$
where $\mu_{1}$ is on the interval between $F_{1}$ and $\mathbb{E}[F_{1}]$ and $\mu_{0}$ is on the interval between $1/J\sum_{i=2}^{J+1}F_{i}$ and $\mathbb{E}[1/J\sum_{i=2}^{J+1}F_{i}]$, and hence, by Assumption 4, we have $l_{f}\leq\mu_{0},\mu_{1}\leq L_{f}$. Therefore,
$\left|\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}]}-\mathbb{E}\left[\frac{F_{1}}{\frac{1}{J}\sum_{i=2}^{J+1}F_{i}}\right]\right|\leq\frac{L_{f}~{}var\left(\frac{1}{J}\sum_{i=2}^{J+1}F_{i}\right)}{l_{f}^{3}}-\frac{cov\left(F_{1},\frac{1}{J}\sum_{i=2}^{J+1}F_{i}\right)}{l_{f}^{2}}.$
Note that by Assumption 4,
$-cov\left(F_{1},\frac{1}{J}\sum_{i=2}^{J+1}F_{i}\right)\leq\left|\frac{1}{J}\sum_{i=2}^{J+1}cov(F_{1},F_{i})\right|\leq\tau_{1}(J),$
and
$var\left(\frac{1}{J}\sum_{i=2}^{J+1}F_{i}\right)$ $\leq\left|\frac{1}{J^{2}}\sum_{i=2}^{J+1}var(F_{i})\right|+\left|\frac{1}{J^{2}}\sum_{i\neq j}cov(F_{i},F_{j})\right|$
$\leq\frac{L_{y}^{2}}{J}+\tau_{2}(J).$
Therefore, we have
$\left|\psi_{0}-\mathbb{E}[\hat{\psi}^{eq2}]\right|$ $\leq L_{y}\left|\frac{\mathbb{E}[F_{1}]}{\mathbb{E}[\frac{1}{J}\sum_{i=2}^{J+1}F_{i}]}-\mathbb{E}\left[\frac{F_{1}}{\frac{1}{J}\sum_{i=2}^{J+1}F_{i}}\right]\right|+C(J)$
$\leq L_{y}\left[\frac{l_{f}\tau_{1}(J)+L_{f}\tau_{2}(J)+L_{f}L_{y}^{2}/J}{l_{f}^{3}}\right]+C(J).$
A.4 Proof of Theorem 4
Consider the model presented in Assumption 6.
$Y_{i,s}=\varrho_{s}+\varphi_{s}^{\top}X_{i}+\vartheta_{s}^{\top}\mu_{i}+\alpha_{i,s}A_{i,s}+\varepsilon_{i,s},$
$F_{i,t}=\rho_{t}+\phi_{t}^{\top}Z_{i}+\theta_{t}^{\top}\mu_{i}+\epsilon_{i,t}.$
Note that under the above model, the potential outcome $Y_{i,s}$ under control will be
$Y_{i,s}^{(0)}=\varrho_{s}+\varphi_{s}^{\top}X_{i}+\vartheta_{s}^{\top}\mu_{i}+\varepsilon_{i,s}.$
The weighted average of the potential outcomes of the donors is
$\sum_{i=2}^{J+1}w_{i}Y_{i,s}=\sum_{i=2}^{J+1}w_{i}Y_{i,s}^{(0)}=\varrho_{s}+\varphi_{s}^{\top}\sum_{i=2}^{J+1}w_{i}X_{i}+\vartheta_{s}^{\top}\sum_{i=2}^{J+1}w_{i}\mu_{i}+\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s}.$
Hence, we have
$Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}=\varphi_{s}^{\top}(X_{1}-\sum_{i=2}^{J+1}w_{i}X_{i})+\vartheta_{s}^{\top}(\mu_{1}-\sum_{i=2}^{J+1}w_{i}\mathbf{\mu}_{i})+\varepsilon_{1,s}-\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s}.$ (11)
Let $F_{i}^{R}$ be the $T\times 1$ vector with the $t$-th element equal to $F_{i,t}$. Similarly, let $\epsilon_{i}^{R}$ be the $T\times 1$ vector with the $t$-th element equal to $\epsilon_{i,t}$. Finally, let $\phi^{R}$ and $\theta^{R}$ be the $T\times d_{r}$ matrix and $T\times d_{u}$ matrices with the $t$-th rows equal to $\phi_{t}$ and $\theta_{t}$, respectively. We have
$F_{1}^{R}-\sum_{i=2}^{J+1}w_{i}F_{i}^{R}=\phi^{R}(Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i})+\theta^{R}(\mu_{1}-\sum_{i=2}^{J+1}w_{i}\mu_{i})+\epsilon_{1}^{R}-\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}.$
Based on Assumption 6, the matrix ${\theta^{R}}^{\top}\theta^{R}$ is invertible. Multiplying by $\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}$ yields
$\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{F_{1}^{R}-\sum_{i=2}^{J+1}w_{i}F_{i}^{R}\}=$ $~{}\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\phi^{R}(Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i})$
$+\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\theta^{R}(\mu_{1}-\sum_{i=2}^{J+1}w_{i}\mu_{i})$
$+\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{\epsilon_{1}^{R}-\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}\},$
which implies that
$\vartheta_{s}^{\top}(\mu_{1}-\sum_{i=2}^{J+1}w_{i}\mu_{i})=$ $~{}\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{F_{1}^{R}-\sum_{i=2}^{J+1}w_{i}F_{i}^{R}\}$ (12)
$-\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\phi^{R}(Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i})$
$-\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{\epsilon_{1}^{R}-\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}\}.$
From Equations (11) and (12), we conclude that
$Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}=$ $~{}\varphi_{s}^{\top}(X_{1}-\sum_{i=2}^{J+1}w_{i}X_{i})$
$+\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{F_{1}^{R}-\sum_{i=2}^{J+1}w_{i}F_{i}^{R}\}$
$-\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\phi^{R}(Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i})$
$-\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{\epsilon_{1}^{R}-\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}\}$
$+\varepsilon_{1,s}-\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s}$
$=$ $~{}-\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{\epsilon_{1}^{R}-\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}\}$
$+\varepsilon_{1,s}-\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s},$
where the last equality follows from Conditions (4)-(6).
Let
$R_{1s}=\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}$
$R_{2s}=-\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\epsilon_{1}^{R}$
$R_{3s}=\varepsilon_{1,s}-\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s}.$
Then,
$Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}=R_{1s}+R_{2s}+R_{3s}.$
For $t\in\{1,\dots,T\}$, the noise variable $\epsilon_{i,t}$ has mean zero, which gives us $R_{2s}=0$. $R_{3s}$ also has mean zero as the weights obtained from the reference domain are independent of $\varepsilon_{i,s}$. Only $R_{1s}$ may not have mean zero as the weights are dependent on $\epsilon_{i,t}$. We re-write $R_{1s}$ as
$R_{1s}$ $=\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}$
$=\sum_{i=2}^{J+1}{w_{i}}\sum_{t=1}^{T}\vartheta_{s}^{\top}\{\sum_{l=1}^{T}{\theta_{l}}\theta_{l}^{\top}\}^{-1}\theta_{t}\epsilon_{i,t}.$
We note that
$\{\vartheta_{s}^{\top}\{\sum_{l=1}^{T}\theta_{l}^{\top}\theta_{l}\}^{-1}\theta_{t}\}^{2}$ $\leq\left(\vartheta_{s}^{\top}\left\{\sum_{l=1}^{T}\theta_{l}\theta_{l}^{\top}\right\}^{-1}\vartheta_{s}\right)\left(\theta_{t}^{\top}\left\{\sum_{l=1}^{T}\theta_{l}\theta_{l}^{\top}\right\}^{-1}\theta_{t}\right)$
$\leq\left(\frac{\bar{\vartheta}^{2}d_{u}}{T\underline{\xi}}\right)\left(\frac{\bar{\theta}^{2}d_{u}}{T\underline{\xi}}\right).$
Let
$u_{is}$ $=\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\epsilon_{i}^{R}$
$=\sum_{t=1}^{T}\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta_{t}}\epsilon_{i,t}.$
Therefore, $u_{is}$, which is a linear combination of independent sub-Gaussian with variance proxy $\bar{\sigma}^{2}$, will be sub-Gaussian with variance proxy $\left({\bar{\vartheta}\bar{\theta}d_{u}}/{\underline{\xi}}\right)^{2}\bar{\sigma}^{2}/T$.
Let $\mathbf{W}=\{w\in\mathbb{R}^{J};\sum w_{i}=1\}$. Theorem 1.16 from (Rigollet and Hütter,, 2023) implies
$\left|\mathbb{E}\left[Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}\right]\right|$ $=\left|\mathbb{E}\left[\sum_{i=2}^{J+1}w_{i}u_{is}\right]\right|$
$\leq\mathbb{E}\left[\max_{w\in\mathbf{W}}\left|\sum_{i=2}^{J+1}w_{i}u_{is}\right|\right]$
$\leq\sqrt{2}\bar{\vartheta}\bar{\theta}d_{u}\bar{\sigma}\underline{\xi}^{-1}\sqrt{\frac{\log(2J)}{T}}.$
Hence, we can bound the bias of our proposed estimator as
$\left|\mathbb{E}[\hat{\psi}^{sc}]-\psi_{0}\right|$ $=\left|\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}\left[Y_{1,s}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}\right]-\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}\left[Y_{1,s}-Y_{1,s}^{(0)}\right]\right|$
$=\left|\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}\left[Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}\right]\right|$
$\leq\frac{1}{S}\sum_{s=1}^{S}\left|\mathbb{E}\left[Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}\right]\right|$
$\leq\sqrt{2}\bar{\vartheta}\bar{\theta}d_{u}\bar{\sigma}\underline{\xi}^{-1}\sqrt{\frac{\log(2J)}{T}}.$